A is x years old and B is (5x-12) years old. Given that B is twice as old as A, the the age of A is: ...?
Answer by stanbon(75887)

A is x years old and B is (5x-12) years old. Given that B is twice as old as A, the the age of A is: ...?
-----
Equation:
B = 2*A
-----
5x-12 = 2*x
----
3x = 12
x = 4 yrs (A's age now)
================
Cheers,
Stan H.
---------------